OpenStudy (luckylu):

I need help I am not sure how to get the answer for this problem. Determine the equation of the graph and select the correct answer below.

OpenStudy (luckylu):

Answer Choices y = (x + 3)^2 − 2 y = (x − 3)^2 − 2 y = (x + 3)^2 + 2 y = (x − 3)^2 + 2

OpenStudy (jadeishere):

The parent graph's equation is a quadratic, which is f(x) = x^2 The vertex of the parabola is at (0, 0) for the parent. Now the vertex is at (-3, 2) which means it moved left 3 and up 2 f(x) = a(x - h) + k h = x value k= y value

OpenStudy (luckylu):

so that means (x+3) +2?

OpenStudy (luckylu):

I got it right thank you very much :D

OpenStudy (jadeishere):

Yeah! Great job :)